OpenStudy (anonymous):

36 = y - 5y - 12 someone help me please?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

@bianca214 Are you still there?

OpenStudy (skullpatrol):

Do you want to learn how to do this on your own?

OpenStudy (anonymous):

no

OpenStudy (skullpatrol):

$$36 = y - 5y - 12$$ $$36=(1-5)y-12$$

OpenStudy (skullpatrol):

Can you do the next step?

OpenStudy (anonymous):

yea thanks